# Persona Dynamics: Unveiling the Impact of Personality Traits on Agents in Text-Based Games

### v1

@@ D.6 Detailed Criteria for Place Classification
D.7 World Visualization and Trajectory of Agent
We applied the visualization method presented in Section 5.1 to all agents, with results shown in Figure 13 to 16. Additionally, we visualize the map of the game Zork-1, Zork-2, and Zork-3 from Jiminy Cricket benchmark in in Figure 18 to 20.
D.8 Prompt used in GPT-4.
@@ E.2 BFI and SD-3
To measure these personality traits, psychologists have developed various assessment tools. The Big Five Inventory (BFI) is one of the most commonly used instruments for assessing the Big Five personality traits. It consists of a series of statements that respondents rate based on how accurately they reflect their own behavior and preferences. For assessing the Dark Triad traits, the Short Dark Triad (SD-3) questionnaire is widely used. The SD-3 is a brief yet effective measure, designed to assess Machiavellianism, Narcissism, and Psychopathy with just a few items per trait. Full set of BFI and SD-3 are in listed in Table 21 and  22.

### v2

@@ D.5 Detailed Criteria for Place Classification
D.6 Prompt used for GPT-4
@@
D.7 Human gameplay data & Human annotation
In this work, humans were employed for two tasks. For generating human gameplay trajectory (§6.2), three individuals each played 10 games, resulting in two trajectories for each of the 15 games covered in the main paper. To facilitate gameplay, valid action candidates provided by the environment were made available, and only 100 actions were collected. Each game took approximately 5-10 minutes on average. After gameplay, they proceeded to annotate the personality types of selected actions comprising other players’ trajectories. To ensure objectivity, they evaluated games they hadn’t played themselves. Similar to the classifier introduced in §3, human annotators viewed state and action pairs $(s_{t},a_{t})$ and assessed whether $a_{t}$ exhibited high tendency, low tendency, or neutrality with respect to the given personality type. We pay each annotator 15\$ per hour.
@@ E.2 BFI and SD-3
To measure these personality traits, psychologists have developed various assessment tools. The Big Five Inventory (BFI) is one of the most commonly used instruments for assessing the Big Five personality traits. It consists of a series of statements that respondents rate based on how accurately they reflect their own behavior and preferences. For assessing the Dark Triad traits, the Short Dark Triad (SD-3) questionnaire is widely used. The SD-3 is a brief yet effective measure, designed to assess Machiavellianism, Narcissism, and Psychopathy with just a few items per trait. Full set of BFI and SD-3 are in listed in Table 24 and  25.